Given that u=u(x,y,z)=c1,v=v(x,y,z)=c2 are solutions of dx/P=dy/Q=dz/R. Show that F(u,v)=0 is a general solution of the Lagrange's equation

Pp+Qq=R


1
Expert's answer
2021-11-01T18:48:07-0400

Given Lagrange’s partial differential equation is


"\ud835\udc43\ud835\udc5d + \ud835\udc44\ud835\udc5e = \ud835\udc45"

Let

"F(u,v)=0"

be the solution of the given Lagrange’s equation .

Differentiating partially with respect to "x" , we have


"(\dfrac{\partial F}{\partial u}\dfrac{\partial u}{\partial x}+\dfrac{\partial F}{\partial u}\dfrac{\partial u}{\partial z}\dfrac{\partial z}{\partial x})"

"+(\dfrac{\partial F}{\partial v}\dfrac{\partial v}{\partial x}+\dfrac{\partial F}{\partial v}\dfrac{\partial v}{\partial z}\dfrac{\partial z}{\partial x})=0"

Or


"\dfrac{\partial F}{\partial u}(\dfrac{\partial u}{\partial x}+p\dfrac{\partial u}{\partial z})+\dfrac{\partial F}{\partial v}(\dfrac{\partial v}{\partial x}+p\dfrac{\partial v}{\partial z})=0"



Similary, differentiating partially with respect to "x" , we have


"\dfrac{\partial F}{\partial u}(\dfrac{\partial u}{\partial y}+q\dfrac{\partial u}{\partial z})+\dfrac{\partial F}{\partial v}(\dfrac{\partial v}{\partial y}+q\dfrac{\partial v}{\partial z})=0"


Eliminating "\dfrac{\partial F}{\partial u}" and "\dfrac{\partial F}{\partial v}," we get


"(\dfrac{\partial u}{\partial x}+p\dfrac{\partial u}{\partial z})(\dfrac{\partial v}{\partial y}+q\dfrac{\partial v}{\partial z})"

"-(\dfrac{\partial u}{\partial y}+q\dfrac{\partial u}{\partial z})(\dfrac{\partial v}{\partial x}+p\dfrac{\partial v}{\partial z})=0"

Then

"(\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial z}-\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial y})p+(\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial x}-\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial z})q"

"=\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial y}-\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial x}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (*)"

which can also be put in the form


"\dfrac{\partial (u,v)}{\partial (y,z)}p+\dfrac{\partial (u,v)}{\partial (z,x)}q=\dfrac{\partial (u,v)}{\partial (x,y)}"

or


"Pp+Qq=R"

where


"P=\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial z}-\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial y}=\dfrac{\partial (u,v)}{\partial (y,z)}"

"Q=\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial x}-\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial z}=\dfrac{\partial (u,v)}{\partial (z,x)}"

"R=\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial y}-\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial x}=\dfrac{\partial (u,v)}{\partial (x,y)}"

Thus, the equation "\ud835\udc43\ud835\udc5d + \ud835\udc44\ud835\udc5e = \ud835\udc45" is a partial differential equation of order one and degree one for which "F(u,v)" is a solution.


Now taking the differentials of two independent solutions "\ud835\udc62(\ud835\udc65, \ud835\udc66, \ud835\udc67) = \ud835\udc50_1" and "\ud835\udc63(\ud835\udc65, \ud835\udc66, \ud835\udc67) = \ud835\udc50_2 ," we get


"\dfrac{\partial u}{\partial x}dx+\dfrac{\partial u}{\partial y}dy+\dfrac{\partial u}{\partial z}dz=0"

"\dfrac{\partial v}{\partial x}du+\dfrac{\partial v}{\partial y}dy+\dfrac{\partial v}{\partial z}dz=0"

Since "u" and "v" are independent functions, therefore, solving equations for the ratios "\ud835\udc51\ud835\udc65, \ud835\udc51\ud835\udc66,\ud835\udc51\ud835\udc67," we get


"\dfrac{dx}{\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial z}-\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial y}}=\dfrac{dy}{\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial x}-\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial z}}"

"=\dfrac{dz}{\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial y}-\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial x}}"

Given that "u=u(x,y,z)=c_1,v=v(x,y,z)=c_2" are solutions of


"dx\/P=dy\/Q=dz\/R"

Then


"\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial z}-\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial y}=kP"

"\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial x}-\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial z}=kQ"

"\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial y}-\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial x}=kR"

Substitute in"(*)"


"kPp+kQq=kR"

We get


"Pp+Qq=R"

which is the given partial differential equation.

Therefore, if "\ud835\udc62(\ud835\udc65, \ud835\udc66, \ud835\udc67) = \ud835\udc50_1" and "\ud835\udc63(\ud835\udc65, \ud835\udc66, \ud835\udc67) = \ud835\udc50_2" are two independent solutions of the system of differential equations "dx\/P=dy\/Q=dz\/R," then "F(u,v)=0" is a solution of "\ud835\udc43\ud835\udc5d + \ud835\udc44\ud835\udc5e = \ud835\udc45," where "F" is an arbitrary function.